a 10.0L cylinder of gas is stored at room temperature (20.0c) and a pressure of 1800psi if the gas is transferred to a 6.0L cylinder at what temperature would it have to be stored in order for the pressure to remain at 1800 psi


1
Expert's answer
2022-05-20T05:09:04-0400

Given:

V1 = 10.0 L

T1 = 20.0°C + 273 = 293 K

P1 = P2 = 1800 psi = constant

V2 = 6.0 L

T2 = unknown


Formula: V1/T1 = V2/T2


Solution:

Since the pressure and amount of gas remain unchanged, Charles's law can be used.

Charles's law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the resulting temperature, solve the equation for T2:

T2 = V2T1 / V1

T2 = (6.0 L × 293 K) / (10.0 L) = 175.8 K = 176 K

T2 = 176 K (or −97.0°C)


Answer: at −97.0°C